A number of two digits is increased by 9. When the digits are reversed, the sum of the digits is 5. Find the number?
1
Expert's answer
2020-06-02T17:53:53-0400

The original condition is:

The sum of the digits of a number lying between 10 and 100 is 5. When the digits are reversed, the number is increased by 9. Find the number.

Assume that the initial number is


"10a+b."

If the digits are reversed, the number of two digits is increased by 9:


"10b+a=10a+b+9,\\
9b-9a=9,\\
b-a=1."

The sum of the digits is 5:


"a+b=5."

Therefore, we have two simple equation that form a system. The solution of the system


"b-a=1\\
a+b=5"

is


"a=2\\
b=3."